# Application of AI to formal methods — an analysis of current trends

### v1

@@ 8 Related work
While the lack of available literature partially motivated this work, there are noteworthy related works. One of the first contributions one finds when investigating the topic is the one of Amrani et al. [52]. Here, the authors came to a similar conclusion as we did about the complicated nature of the investigated landscape of available literature. Therefore, they also resorted to a complex search string. The evaluation and focus, however, are different. The authors limit themselves to an excerpt of initial results and only conduct backward snowballing, leaving room for whether the corpus is complete. Additionally, the research questions are more specifically tailored to individual FM topics, while our work aims to give a general overview.
Wang et al. [53] aims to give a general taxonomy of learning-based FM. One core contribution is that they divide the topic into two general directions: the learning of formal specifications and learning for formal verification. They proceeded to structure the literature found by individuals underneath this taxonomy. Compared to their work, our work aims to give a general introduction to the cross-section of the two fields while the authors dive into the specifics of learning algorithms for formal verification.
@@
During our search, we found many contributions that targeted the general constraint satisfaction problem. However, we excluded these results as they are only loosely connected with FM. Amadini et al. [61] conducted a general evaluation of portfolio approaches for CSP problems. Popescu et al. [62] gives an overview of machine learning techniques for CSP problems. While Yang et al. [63] conducts a detailed survey of intelligent optimization algorithms for satisfaction problems.
@@
Multiple works aim to overview the topic within the math and automated reasoning community. Urban and Vyskočil [64] give an overview of AI-supported automated theorem proving. Blanchette et al. [65] surveyed ML-supported axiom selection, and two years later, Blanchette et al. [66] surveyed the rising interest in automatizing reasoning over proofs. England [67] published a survey on ML for symbolic reasoning. Tran et al. [68] gave a short overview of the emerging field of NLP for premise selection.
@@
Brunello et al. [72] and Buzhinsky [73] published surveys on creating temporal logic from natural language, while Fuggitti and Chakraborti [74]. Szegedy [75] discusses the topic of auto formalization of natural language requirements.
@@
Less prominent were the topics of model repair, where Barriga et al. [76] published a longer article about the state of the art. Haltermann and Wehrheim [77], gives an overview of invariant generation. Pan and Mishra [78] surveyed hardware vulnerability analysis via ML, a field to which we found no contributions. For model checking, Besbas et al. [79] gave a brief literature review of four pages.
@@ 9 Conclusion
In this systematic mapping study, we present the results of our work on assessing the quantity of research in applying artificial intelligence to formal methods. Overall, we found 457, from which we selected 189 for a closer investigation. Concluding from this investigation, we see the current trend is yet to mature, as many contributions are making some practical application. However, only a few studies aim to create theoretical groundwork, benchmarks, or case studies. Furthermore, we see a focus on theorem proving and SAT solving. Model checking and model synthesis are underrepresented compared to this. Later may be surprising with the advantage of LLM that becomes precise in creating output after a user’s desire. Most work uses some neural nets, reinforcement learning, or a combination of multiple approaches.
Appendix A Tables for RQ 1
Entries grouped by years
Year Articles Count
@@ Appendix A Tables for RQ 1
Year Articles Count
@@
2020 [112, 113, 114, 115, 116, 117, 118, 119, 120, 121, 122, 123, 124, 125, 126, 127, 128, 129, 130, 131, 132, 133, 134, 135, 136, 137, 138, 139, 140, 141, 142, 143, 144, 145, 146, 147, 148, 149, 150, 151, 152, 153, 154, 155, 156, 157, 158, 159] 48
2021 [160, 161, 162, 163, 164, 165, 166, 167, 168, 169, 170, 171, 172, 173, 174, 175, 176, 177, 178, 179, 180, 181, 182, 183, 184, 185, 186, 187, 188, 189, 190, 191, 192, 193, 194, 195, 196, 197, 198] 39
2022 [199, 200, 201, 202, 203, 204, 205, 206, 207, 208, 209, 210, 211, 212, 213, 214, 215, 216, 217, 218, 219, 220, 221, 222, 223, 224, 225, 226, 227, 228, 229, 230, 231, 232, 233, 234, 235, 236, 237, 238, 239, 240, 241, 242, 243, 244, 245, 246, 247, 248, 249] 51
2023 [250, 251, 252, 253, 254, 255, 256, 257, 258, 259, 260, 261, 262, 60, 263, 264, 265, 266, 267, 268, 269] 21
Entries grouped by type of publication venue
Entry type Articles Count
[…]
Entry type Articles Count
[…]
Book chapter [233] 1
Workshop/short paper [94, 95, 96, 99, 105, 120, 132, 137, 138, 148, 154, 156, 160, 162, 166, 173, 175, 176, 192, 200, 214, 223, 228, 236, 238, 246, 259, 264, 268] 29
Conference paper [81, 83, 84, 85, 88, 89, 90, 91, 92, 93, 97, 101, 102, 103, 104, 106, 107, 113, 115, 116, 117, 118, 119, 121, 123, 124, 125, 126, 127, 128, 129, 130, 134, 139, 140, 141, 142, 144, 145, 146, 149, 150, 151, 152, 153, 155, 157, 158, 161, 163, 164, 167, 168, 169, 171, 172, 174, 177, 178, 179, 181, 182, 183, 184, 185, 186, 189, 190, 191, 193, 194, 195, 196, 197, 198, 201, 202, 203, 204, 205, 206, 207, 208, 209, 210, 211, 212, 216, 217, 218, 220, 221, 222, 224, 225, 227, 229, 230, 231, 232, 234, 235, 239, 240, 242, 243, 244, 245, 247, 252, 254, 256, 260, 261, 263, 265, 266, 267, 269, 110, 111, 159] 122
Entries grouped by type of contribution
Contribution type Articles Count
[…]
Contribution type Articles Count
[…]
Tool improvement [84, 85, 91, 144, 178, 179, 189, 208, 224, 254] 10
Idea [242] 1
Approach [107, 116] 2
Framework [167, 193, 203, 217, 219, 247] 6
Methodology [81, 82, 86, 87, 88, 93, 94, 95, 96, 98, 99, 101, 103, 105, 106, 108, 109, 113, 114, 117, 118, 119, 121, 123, 124, 127, 128, 129, 133, 134, 135, 136, 137, 140, 141, 142, 143, 147, 148, 149, 151, 154, 155, 156, 157, 158, 160, 161, 162, 163, 165, 166, 168, 169, 170, 173, 175, 176, 177, 181, 182, 184, 185, 186, 187, 188, 194, 196, 197, 198, 199, 200, 201, 202, 204, 205, 206, 207, 212, 213, 214, 215, 216, 218, 220, 221, 223, 225, 226, 227, 229, 231, 234, 235, 236, 237, 238, 239, 240, 243, 245, 248, 249, 250, 251, 252, 255, 257, 259, 260, 261, 262, 60, 264, 266, 268, 269, 111] 118
Case study [233, 253] 2
Benchmark [112, 150] 2
Training data [92, 145, 146, 241, 246] 5
Mult. contributions [80, 89, 90, 97, 100, 104, 125, 126, 130, 132, 138, 139, 171, 172, 174, 180, 183, 195, 209, 211, 222, 228, 232, 244, 256, 258, 267, 110, 159] 29
Entries grouped by type of contribution and publication venue
Contribution type Entry type Articles Count
[…]
Contribution type Entry type Articles Count
[…]
Workshop/short paper [120, 192] 2
Conference paper [83, 102, 115, 152, 164, 190, 191, 210, 230, 263] 10
Tool improvement Conference paper [84, 85, 91, 144, 178, 179, 189, 208, 224, 254] 10
Idea Conference paper [242] 1
Approach Conference paper [107, 116] 2
Framework Journal paper [219] 1
Conference paper [167, 193, 203, 217, 247] 5
Methodology Journal paper [82, 86, 87, 98, 108, 109, 114, 133, 135, 136, 143, 147, 165, 170, 187, 188, 199, 213, 215, 226, 237, 248, 249, 250, 251, 255, 257, 262, 60] 29
Workshop/short paper [94, 95, 96, 99, 105, 137, 148, 154, 156, 160, 162, 166, 173, 175, 176, 200, 214, 223, 236, 238, 259, 264, 268] 23
Conference paper [81, 88, 93, 101, 103, 106, 113, 117, 118, 119, 121, 123, 124, 127, 128, 129, 134, 140, 141, 142, 149, 151, 155, 157, 158, 161, 163, 168, 169, 177, 181, 182, 184, 185, 186, 194, 196, 197, 198, 201, 202, 204, 205, 206, 207, 212, 216, 218, 220, 221, 225, 227, 229, 231, 234, 235, 239, 240, 243, 245, 252, 260, 261, 266, 269, 111] 66
Case study Journal paper [253] 1
Book chapter [233] 1
Benchmark Journal paper [112] 1
Conference paper [150] 1
Training data Journal paper [241] 1
Workshop/short paper [246] 1
Conference paper [92, 145, 146] 3
Mult. contributions Journal paper [80, 100, 180, 258] 4
Workshop/short paper [132, 138, 228] 3
Conference paper [89, 90, 97, 104, 125, 126, 130, 139, 171, 172, 174, 183, 195, 209, 211, 222, 232, 244, 256, 267, 110, 159] 22
Appendix B Tables for RQ 2
Entries grouped by type of AI type used
AI technique Articles Count
[…]
AI technique Articles Count
[…]
RF [117, 201] 2
SVM [133] 1
KNN [122, 125] 2
NN [83, 90, 92, 96, 97, 99, 102, 104, 106, 120, 121, 123, 124, 126, 129, 134, 135, 137, 139, 140, 141, 146, 150, 151, 152, 155, 158, 163, 165, 169, 170, 174, 175, 177, 178, 179, 182, 190, 192, 193, 195, 196, 203, 205, 206, 209, 210, 211, 218, 219, 220, 222, 227, 228, 231, 232, 233, 234, 237, 239, 242, 254, 257, 258, 259, 262, 263, 268, 110, 159] 70
RL [94, 98, 100, 101, 115, 118, 127, 130, 132, 142, 148, 149, 161, 167, 168, 171, 176, 184, 191, 199, 207, 224, 230, 235, 240, 241, 250, 253, 256, 260, 264, 267] 32
EA [81, 107, 108, 109, 147, 188, 194, 198, 200, 204, 213, 243, 248, 269] 14
NLP [116, 131, 145, 160, 162, 183, 186, 197, 202, 217, 223, 229, 236, 238, 245, 246] 16
Automaton learning [138, 249] 2
Bayesian inference [113, 136, 221] 3
Clustering [86, 114, 244] 3
Data mining [88, 143] 2
Mult. AI techniques [80, 82, 85, 87, 89, 93, 105, 112, 119, 128, 156, 164, 166, 172, 173, 180, 187, 189, 208, 212, 215, 226, 251, 252, 261, 266, 111] 27
Custom [103, 154, 157, 185, 214, 247, 255, 60] 8
Entries that use multiple AI types
AI technique Articles Count
[…]
AI technique Articles Count
[…]
SVM [85, 93, 166, 180, 215, 226] 6
KNN [85, 87, 166, 172, 226] 5
NN [80, 89, 112, 128, 173, 180, 187, 189, 208, 212, 251, 252, 111] 13
EA [187] 1
Clustering [105, 156, 212] 3
Data mining [119] 1
Logistic models [80, 87, 93, 112, 215, 226] 6
LR [85, 87, 166] 3
[82] 1
Ridge regression [166, 226, 251] 3
Naive bayes [93, 112, 215, 226] 4
Trees [80, 82, 85, 105, 112, 156, 164, 166, 172, 189, 226, 111] 12
Nearest centroid [226] 1
Passive aggressive [226] 1
Stochastic gradient [226] 1
LLM [261, 266] 2
Bagging [85] 1
Adaboost [85, 164] 2
Gaussian process [226] 1
MCTS [261] 1
SMAC [105] 1
Entries grouped by AI types and contribution type
Contribution type AI technique Articles Count
[…]
Contribution type AI technique Articles Count
[…]
NN [83, 102, 120, 152, 190, 192, 210, 263] 8
RL [115, 191, 230] 3
NLP [131] 1
Mult. AI techniques [164] 1
Tool improvement DT [84, 91, 144] 3
NN [178, 179, 254] 3
RL [224] 1
Mult. AI techniques [85, 189, 208] 3
Idea NN [242] 1
Approach EA [107] 1
NLP [116] 1
Framework NN [193, 203, 219] 3
RL [167] 1
NLP [217] 1
Custom [247] 1
Methodology DT [95, 181, 216, 225] 4
RF [117, 201] 2
SVM [133] 1
NN [96, 99, 106, 121, 123, 124, 129, 134, 135, 137, 140, 141, 151, 155, 158, 163, 165, 169, 170, 175, 177, 182, 196, 205, 206, 218, 220, 227, 231, 234, 237, 239, 257, 259, 262, 268] 36
RL [94, 98, 101, 118, 127, 142, 148, 149, 161, 168, 176, 184, 199, 207, 235, 240, 250, 260, 264] 19
EA [81, 108, 109, 147, 188, 194, 198, 200, 204, 213, 243, 248, 269] 13
NLP [160, 162, 186, 197, 202, 223, 229, 236, 238, 245] 10
Automaton learning [249] 1
Bayesian inference [113, 136, 221] 3
Clustering [86, 114] 2
Data mining [88, 143] 2
Mult. AI techniques [82, 87, 93, 105, 119, 128, 156, 166, 173, 187, 212, 215, 226, 251, 252, 261, 266, 111] 18
Custom [103, 154, 157, 185, 214, 255, 60] 7
Case study NN [233] 1
RL [253] 1
Benchmark NN [150] 1
Mult. AI techniques [112] 1
Training data NN [92, 146] 2
RL [241] 1
NLP [145, 246] 2
Mult. contributions KNN [125] 1
NN [90, 97, 104, 126, 139, 174, 195, 209, 211, 222, 228, 232, 258, 110, 159] 15
RL [100, 130, 132, 171, 256, 267] 6
NLP [183] 1
Automaton learning [138] 1
Clustering [244] 1
Mult. AI techniques [80, 89, 172, 180] 4
Entries grouped by AI types and publication years
AI technique Year Articles Count
[…]
AI technique Year Articles Count
[…]
2020 [144] 1
2021 [181] 1
2022 [216, 225] 2
RF 2020 [117] 1
2022 [201] 1
SVM 2020 [133] 1
KNN 2020 [122, 125] 2
NN 2019 [83, 90, 92, 96, 97, 99, 102, 104, 106, 110] 10
2020 [120, 121, 123, 124, 126, 129, 134, 135, 137, 139, 140, 141, 146, 150, 151, 152, 155, 158, 159] 19
2021 [163, 165, 169, 170, 174, 175, 177, 178, 179, 182, 190, 192, 193, 195, 196] 15
2022 [203, 205, 206, 209, 210, 211, 218, 219, 220, 222, 227, 228, 231, 232, 233, 234, 237, 239, 242] 19
2023 [254, 257, 258, 259, 262, 263, 268] 7
RL 2019 [94, 98, 100, 101] 4
2020 [115, 118, 127, 130, 132, 142, 148, 149] 8
2021 [161, 167, 168, 171, 176, 184, 191] 7
2022 [199, 207, 224, 230, 235, 240, 241] 7
2023 [250, 253, 256, 260, 264, 267] 6
EA 2019 [81, 107, 108, 109] 4
2020 [147] 1
2021 [188, 194, 198] 3
2022 [200, 204, 213, 243, 248] 5
2023 [269] 1
NLP 2020 [116, 131, 145] 3
2021 [160, 162, 183, 186, 197] 5
2022 [202, 217, 223, 229, 236, 238, 245, 246] 8
Automaton learning 2020 [138] 1
2022 [249] 1
Bayesian inference 2020 [113, 136] 2
2022 [221] 1
Clustering 2019 [86] 1
2020 [114] 1
2022 [244] 1
Data mining 2019 [88] 1
2020 [143] 1
Mult. AI techniques 2019 [80, 82, 85, 87, 89, 93, 105, 111] 8
2020 [112, 119, 128, 156] 4
2021 [164, 166, 172, 173, 180, 187, 189] 7
2022 [208, 212, 215, 226] 4
2023 [251, 252, 261, 266] 4
Custom 2019 [103] 1
2020 [154, 157] 2
2021 [185] 1
2022 [214, 247] 2
2023 [255, 60] 2
Entries grouped by FM types
FM technique Articles Count
[…]
FM technique Articles Count
[…]
SMT [89, 95, 119, 161, 163, 164, 171, 185, 186, 216, 251, 263, 264] 13
TP [83, 84, 87, 88, 91, 92, 94, 96, 97, 99, 102, 105, 106, 107, 115, 120, 122, 123, 124, 125, 126, 129, 130, 131, 132, 136, 138, 139, 143, 144, 145, 146, 148, 149, 152, 154, 155, 156, 165, 167, 168, 172, 173, 174, 175, 176, 177, 178, 179, 180, 183, 184, 187, 188, 189, 190, 191, 196, 197, 207, 208, 209, 214, 217, 218, 221, 223, 224, 225, 228, 229, 232, 236, 237, 238, 241, 245, 250, 253, 254, 257, 258, 261, 266, 111] 85
MC [81, 82, 108, 109, 113, 118, 166, 195, 198, 205, 220, 226, 230, 248, 249, 252, 268, 269] 18
Synthesis [80, 93, 116, 157, 160, 162, 192, 193, 199, 202, 210, 211, 215, 244, 246, 247, 256, 260, 267] 19
Other [100, 134, 135, 140, 182, 227, 231, 233, 243] 9
Entries grouped by TP application types
Application goal Articles Count
[…]
Application goal Articles Count
[…]
Premise selection [84, 91, 96, 124, 139, 154, 155, 156, 168, 173, 174, 177, 178, 188, 189, 196, 197, 209, 214, 217, 218, 221, 223, 237, 241, 254, 111] 27
Proof search [83, 92, 94, 115, 120, 122, 123, 125, 130, 136, 144, 145, 146, 165, 176, 187, 191, 224, 225, 250, 261] 21
@@ Appendix B Tables for RQ 2
Mult. Techniques [102, 129, 167, 245, 266] 5
Entries grouped by synthesis application types
Application goal Articles Count
[…]
Application goal Articles Count
[…]
@@
Entries grouped by grade of TP automatization and TP application types
Application goal Automation grade Articles Count
[…]
Application goal Automation grade Articles Count
[…]
Premise selection Interactive [91] 1
Automatic [84, 96, 124, 139, 154, 155, 156, 168, 173, 174, 177, 178, 188, 189, 196, 197, 209, 214, 217, 218, 221, 223, 237, 241, 254, 111] 26
Proof search Automatic [83, 92, 94, 115, 120, 122, 123, 125, 130, 136, 144, 145, 146, 165, 176, 187, 191, 224, 225, 250, 261] 21
Proof mining Interactive [143] 1
Automatic [183] 1
Proof synthesis Automatic [97, 99, 126, 131, 138, 149, 175, 190, 207, 208, 228, 229, 232, 236, 238, 253, 257, 258] 18
Symbolic Automatic [184] 1
Formula synthesis Automatic [148] 1
Formula classification Interactive [179] 1
Automatic [180] 1
Tactic prediction Interactive [88, 132, 172] 3
Automatic [106] 1
Lemma name generation Interactive [152] 1
Portfolio solving Automatic [87] 1
Mult. Techniques Interactive [102, 167, 266] 3
Automatic [129] 1
Mixed [245] 1
Entries grouped by level of logic tackled by TP and TP application types
Application goal Formalism Articles Count
[…]
Application goal Formalism Articles Count
[…]
HOL [107] 1
Premise selection FOL [84, 91, 96, 124, 154, 156, 168, 173, 174, 177, 178, 189, 196, 209, 214, 221, 237, 241, 254, 111] 20
HOL [139, 155, 188, 197, 217, 218, 223] 7
Proof search FOL [83, 115, 136, 144, 145, 146, 165, 250] 8
HOL [92, 94, 120, 122, 123, 125, 130, 176, 187, 191, 224, 225, 261] 13
Proof mining HOL [143, 183] 2
Proof synthesis FOL [99, 190, 207, 208, 229, 236, 238] 7
HOL [97, 126, 131, 138, 149, 175, 228, 232, 253, 257, 258] 11
Symbolic HOL [184] 1
Formula synthesis HOL [148] 1
Formula classification FOL [179, 180] 2
Tactic prediction HOL [88, 106, 132, 172] 4
Lemma name generation HOL [152] 1
Portfolio solving HOL [87] 1
Mult. Techniques HOL [102, 129, 167, 245, 266] 5
Entries grouped by FM techniques and years
FM technique Year Articles Count
[…]
FM technique Year Articles Count
[…]
2020 [112, 114, 117, 121, 127, 128, 133, 137, 141, 142, 147, 150, 151, 158, 159] 15
2021 [169, 170, 181, 194] 4
2022 [200, 201, 203, 204, 206, 212, 213, 219, 222, 234, 235, 239, 240, 242] 14
2023 [255, 259, 262, 60] 4
SMT 2019 [89, 95] 2
2020 [119] 1
2021 [161, 163, 164, 171, 185, 186] 6
2022 [216] 1
2023 [251, 263, 264] 3
TP 2019 [83, 84, 87, 88, 91, 92, 94, 96, 97, 99, 102, 105, 106, 107, 111] 15
2020 [115, 120, 122, 123, 124, 125, 126, 129, 130, 131, 132, 136, 138, 139, 143, 144, 145, 146, 148, 149, 152, 154, 155, 156] 24
2021 [165, 167, 168, 172, 173, 174, 175, 176, 177, 178, 179, 180, 183, 184, 187, 188, 189, 190, 191, 196, 197] 21
2022 [207, 208, 209, 214, 217, 218, 221, 223, 224, 225, 228, 229, 232, 236, 237, 238, 241, 245] 18
2023 [250, 253, 254, 257, 258, 261, 266] 7
MC 2019 [81, 82, 108, 109] 4
2020 [113, 118] 2
2021 [166, 195, 198] 3
2022 [205, 220, 226, 230, 248, 249] 6
2023 [252, 268, 269] 3
Synthesis 2019 [80, 93] 2
2020 [116, 157] 2
2021 [160, 162, 192, 193] 4
2022 [199, 202, 210, 211, 215, 244, 246, 247] 8
2023 [256, 260, 267] 3
Other 2019 [100] 1
2020 [134, 135, 140] 3
2021 [182] 1
2022 [227, 231, 233, 243] 4
Entries grouped by FM techniques and contributions types
FM technique Entry type Articles Count
[…]
FM technique Entry type Articles Count
[…]
Workshop/short paper [137, 200, 259] 3
Conference paper [85, 90, 101, 103, 104, 117, 121, 127, 128, 141, 142, 150, 151, 158, 169, 181, 194, 201, 203, 204, 206, 212, 222, 234, 235, 239, 240, 242, 110, 159] 30
SMT Journal paper [251] 1
Workshop/short paper [95, 264] 2
Conference paper [89, 119, 161, 163, 164, 171, 185, 186, 216, 263] 10
TP Journal paper [87, 122, 131, 136, 143, 165, 180, 187, 188, 237, 241, 250, 253, 257, 258] 15
Workshop/short paper [94, 96, 99, 105, 120, 132, 138, 148, 154, 156, 173, 175, 176, 214, 223, 228, 236, 238] 18
Conference paper [83, 84, 88, 91, 92, 97, 102, 106, 107, 115, 123, 124, 125, 126, 129, 130, 139, 144, 145, 146, 149, 152, 155, 167, 168, 172, 174, 177, 178, 179, 183, 184, 189, 190, 191, 196, 197, 207, 208, 209, 217, 218, 221, 224, 225, 229, 232, 245, 254, 261, 266, 111] 52
MC Journal paper [82, 108, 109, 226, 248, 249] 6
Workshop/short paper [166, 268] 2
Conference paper [81, 113, 118, 195, 198, 205, 220, 230, 252, 269] 10
Synthesis Journal paper [80, 199, 215] 3
Workshop/short paper [160, 162, 192, 246] 4
Conference paper [93, 116, 157, 193, 202, 210, 211, 244, 247, 256, 260, 267] 12
Other Journal paper [100, 135] 2
Book chapter [233] 1
Conference paper [134, 140, 182, 227, 231, 243] 6
Entries grouped by FM and AI techniques
FM technique AI technique Articles Count
[…]
FM technique AI technique Articles Count
[…]
RF [117, 201] 2
SVM [133] 1
NN [90, 104, 121, 137, 141, 150, 151, 158, 169, 170, 203, 206, 219, 222, 234, 239, 242, 259, 262, 110, 159] 21
RL [98, 101, 127, 142, 235, 240] 6
EA [147, 194, 200, 204, 213] 5
Clustering [86, 114] 2
Mult. AI techniques [85, 112, 128, 212] 4
Custom [103, 255, 60] 3
SMT DT [95, 216] 2
NN [163, 263] 2
RL [161, 171, 264] 3
NLP [186] 1
Mult. AI techniques [89, 119, 164, 251] 4
Custom [185] 1
TP DT [84, 91, 144, 225] 4
KNN [122, 125] 2
NN [83, 92, 96, 97, 99, 102, 106, 120, 123, 124, 126, 129, 139, 146, 152, 155, 165, 174, 175, 177, 178, 179, 190, 196, 209, 218, 228, 232, 237, 254, 257, 258] 32
RL [94, 115, 130, 132, 148, 149, 167, 168, 176, 184, 191, 207, 224, 241, 250, 253] 16
EA [107, 188] 2
NLP [131, 145, 183, 197, 217, 223, 229, 236, 238, 245] 10
Automaton learning [138] 1
Bayesian inference [136, 221] 2
Data mining [88, 143] 2
Mult. AI techniques [87, 105, 156, 172, 173, 180, 187, 189, 208, 261, 266, 111] 12
Custom [154, 214] 2
MC NN [195, 205, 220, 268] 4
RL [118, 230] 2
EA [81, 108, 109, 198, 248, 269] 6
Automaton learning [249] 1
Bayesian inference [113] 1
Mult. AI techniques [82, 166, 226, 252] 4
Synthesis NN [192, 193, 210, 211] 4
RL [199, 256, 260, 267] 4
NLP [116, 160, 162, 202, 246] 5
Clustering [244] 1
Mult. AI techniques [80, 93, 215] 3
Custom [157, 247] 2
Other NN [134, 135, 140, 182, 227, 231, 233] 7
RL [100] 1
EA [243] 1
Entries grouped by TP and AI techniques
Application goal AI technique Articles Count
[…]
Application goal AI technique Articles Count
[…]
Mult. AI techniques [105] 1
Premise selection DT [84, 91] 2
NN [96, 124, 139, 155, 174, 177, 178, 196, 209, 218, 237, 254] 12
RL [168, 241] 2
EA [188] 1
NLP [197, 217, 223] 3
Bayesian inference [221] 1
Mult. AI techniques [156, 173, 189, 111] 4
Custom [154, 214] 2
Proof search DT [144, 225] 2
KNN [122, 125] 2
NN [83, 92, 120, 123, 146, 165] 6
RL [94, 115, 130, 176, 191, 224, 250] 7
NLP [145] 1
Bayesian inference [136] 1
Mult. AI techniques [187, 261] 2
Proof mining NLP [183] 1
Data mining [143] 1
Proof synthesis NN [97, 99, 126, 175, 190, 228, 232, 257, 258] 9
RL [149, 207, 253] 3
NLP [131, 229, 236, 238] 4
Automaton learning [138] 1
Mult. AI techniques [208] 1
Symbolic RL [184] 1
Formula synthesis RL [148] 1
Formula classification NN [179] 1
Mult. AI techniques [180] 1
Tactic prediction NN [106] 1
RL [132] 1
Data mining [88] 1
Mult. AI techniques [172] 1
Lemma name generation NN [152] 1
Portfolio solving Mult. AI techniques [87] 1
Mult. Techniques NN [102, 129] 2
RL [167] 1
NLP [245] 1
Mult. AI techniques [266] 1
Entries grouped by SAT and AI techniques
Application goal AI technique Articles Count
[…]
Application goal AI technique Articles Count
[…]
NN [104, 203, 242, 159] 4
RL [101, 235] 2
Mult. AI techniques [112] 1
Custom [103] 1
Predict solvability DT [181] 1
NN [90, 121, 141, 219] 4
MaxSat NN [150, 170] 2
RL [98, 240] 2
EA [194, 200, 204, 213] 4
Custom [255] 1
SAT parameter selection RF [201] 1
Branching heuristics RL [127] 1
Mult. AI techniques [128] 1
Algorithm/solver selection RF [117] 1
SVM [133] 1
RL [142, 235] 2
Mult. AI techniques [85, 212] 2
Custom [60] 1
Instance selection NN [137, 158, 169, 262] 4
Clustering [86, 114] 2
Dependency analysis NN [259] 1
Meta analysis/performance prediction NN [151, 206, 234] 3
EA [147] 1
Generation NN [239, 110] 2
Mult. SAT tech NN [222] 1
EA [194] 1
Entries grouped by synthesis and AI techniques
Application goal AI technique Articles Count
[…]
Application goal AI technique Articles Count
[…]
RL [256, 260] 2
Loop invariant learning NN [192] 1
RL [267] 1
Clustering [244] 1
Custom [157] 1
Model repair Mult. AI techniques [80, 93, 215] 3
Models/Specification NN [210, 211] 2
RL [199] 1
NLP [116, 160, 202, 246] 4
Custom [247] 1
Annotations NLP [162] 1

### v2

@@ 8 Related work
While the lack of available literature partially motivated this work, there are noteworthy related works. One of the first contributions one finds when investigating the topic is that of Amrani et al. [51]. Here, the authors came to a similar conclusion as we did about the complicated nature of the investigated landscape of available literature. Therefore, they also resorted to a complex search string. The evaluation and focus, however, are different. The authors limit themselves to an excerpt of initial results and only conduct backward snowballing, leaving room for whether the corpus is complete. Additionally, the research questions are more specifically tailored to individual FM topics, while our work aims to give a general overview.
Wang et al. [52] aims to give a general taxonomy of learning-based FM. One core contribution is that they divide the topic into two general directions: the learning of formal specifications and learning for formal verification. They proceeded to structure the literature found by individuals under this taxonomy. Compared to their work, our work aims to give a general introduction to the cross-section of the two fields, while the authors dive into the specifics of learning algorithms for formal verification.
@@
We found many contributions that targeted the general constraint satisfaction problem during our search. However, we excluded these results as they are only loosely connected with FM@. Amadini et al. [60] conducted a general evaluation of portfolio approaches for CSP problems. Popescu et al. [61] gives an overview of machine learning techniques for CSP problems.
@@
Multiple works aim to provide an overview of the topic within the math and automated reasoning community. Urban and Vyskočil [62] give an overview of AI-supported automated theorem proving. Blanchette et al. [63] surveyed ML-supported axiom selection, and two years later, Blanchette et al. [64] surveyed the rising interest in automatizing reasoning over proofs. England [65] published a survey on ML for symbolic reasoning. Tran et al. [66] gave a short overview of the emerging field of NLP for premise selection.
@@
Brunello et al. [70] and Buzhinsky [71] published surveys on creating temporal logic from natural language, while Fuggitti and Chakraborti [72] and Szegedy [73] discuss the topic of auto-formalization of natural language requirements.
@@
Less prominent were the topics of model repair, where Barriga et al. [74] published a longer article about the state of the art. Haltermann and Wehrheim [75] gives an overview of invariant generation. Pan and Mishra [76] surveyed hardware vulnerability analysis via ML, a field to which we found no contributions. For model checking, Besbas et al. [77] gave a brief literature review of four pages.
@@ 9 Conclusion
In this systematic mapping study, we present the results of our work on assessing the quantity of research in applying artificial intelligence to formal methods. Overall, we found 457, from which we selected 189 for a closer investigation. Concluding from this investigation, we see the current trend is yet to mature, as many contributions are making some practical applications. However, only a few studies aim to create theoretical groundwork, benchmarks, or case studies. Furthermore, we see a focus on theorem proving and SAT solving. Model checking and model synthesis are underrepresented compared to this. Most work uses neural nets, reinforcement learning, or a combination of multiple approaches. Noteworthy is the predominant degradation of AI to a support function within existing formal methods.
@@
Supplementary information
The supplementary material provides an overview of the sorted, investigated primary studies. It consists of tables, the basis for the figures presented in this work.
Appendix A Investigated studies
Investigated primary studies sorted by their FM application
FM technique Application Articles Count
[…]
FM technique Application Articles Count
@@ Appendix A Investigated studies
Predict solvability [88, 89, 90, 91, 92] 5
MaxSat [93, 94, 95, 96, 97, 98, 99, 100, 101] 9
SAT parameter selection [102] 1
Branching heuristics [103, 104] 2
Algorithm/solver selection [80, 59, 105, 82, 106, 100, 107, 108] 8
Instance selection [109, 110, 111, 112, 113, 114] 6
Dependency analysis [115] 1
Meta analysis/performance prediction [116, 117, 118, 119] 4
Generation [120, 121] 2
Model counting [81] 1
SMT Quantifier instantiation [122, 123] 2
Solver selection/scheduling [124, 125, 126, 127, 128, 129, 130, 131] 8
Quality assesment [132, 133, 134] 3
TP Heuristic selection [135, 136] 2
Premise selection [137, 138, 139, 140, 141, 142, 143, 144, 145, 146, 147, 148, 149, 150, 151, 152, 153, 154, 155, 156, 157, 158, 159, 160, 161, 162, 163, 164, 165, 166, 167] 31
Proof search [168, 169, 170, 171, 172, 173, 174, 175, 176, 177, 178, 179, 180, 181, 182, 183, 184, 185, 186, 187, 188, 189] 22
@@
Position prediction [150] 1
Symbol guessing [159] 1
MC [220, 221, 222, 223, 224, 225, 226, 227, 228, 229, 230, 231, 232, 233, 234, 235, 236, 237] 18
Synthesis Invariant learning [238, 239, 240] 3
@@
Other Algorithm selection [257, 258] 2
Termination anlysis [259] 1
Program analysis [260, 261, 262] 3
LTL satisfiability [263] 1
General verification [264, 265] 2
@@
Declarations
[…]
*Funding. The authors have no relevant financial or non-financial interests to disclose.
[…]
*Ethical approval. Not applicable.
[…]
*Informed consent. Not applicable.
[…]
*Author Contributions. The first author conceived the idea of conducting the systematic mapping study. The first and second authors conducted the search process outlined in Section 4 together. The application of IC/EC was done independently by the first and second authors. Disagreements in this filtering process were solved by the third author. The first and second authors were responsible for the manuscript’s writing. The third author reviewed the manuscript, provided feedback, and supervised the project. The public repository containing the dataset of found and investigated studies was prepared by the second author.
[…]
*Data Availability Statement. The primary studies found in this systematic mapping study are publicly available at https://github.com/hhu-stups/ai4fm-studies.
[…]
*Conflict of Interest. The authors have no competing interests to declare that are relevant to the content of this article.
[…]
*Clinical Trial Number. Not applicable.